find the x-intercepts.
Y= x^2+ x - 6

Answer by stanbon(57260)

find the x-intercepts.
Y= x^2+ x - 6
----------------
Let y = 0 and solve for "x":
x^2+x-6=0
Factor to get:
(x+3)(x-2)=0
x = -3 or x=2
=============
Cheers,
Stan H.